# Learning to Infer Generative Template Programs for Visual Concepts

### v1

@@ Abstract
neurosymbolic, visual program, concept learning, program induction
@@ 1 Introduction
We experimentally validate that our method is capable of inferring Template Programs across multiple visual domains: 2D layouts, Omniglot characters, and 3D shapes. We demonstrate that Template Programs natively support a number of downstream applications, including few-shot generation and co-segmentation. We are unaware of any other method that is able to perform these tasks in a domain-general fashion, so we compare against either task-specific or domain-specific alternatives. With respect to task-specific approaches, we find that our neurosymbolic method achieves superior performance. For the one domain, Omniglot (Lake et al., 2015), where task-general methods have been proposed, we compare our domain-general method against domain-specific approaches and find that we are able to achieve competitive performance. Code and data will be publicly released.
@@ 2 Related Work
Concept Learning
Many works have studied concept learning and related tasks (Lake et al., 2019b). A number of methods take a task-specific focus, such as training neural networks to perform few-shot classification (Vinyals et al., 2016; Snell et al., 2017). Other approaches have focused instead on generation, under different conditioning frameworks (Rezende et al., 2016; Giannone et al., 2022). While these methods perform well in the areas they specialize for, none of them achieve the task-flexibility we desire.
@@
Inverse Procedural Modeling
While not often presented as concept learners, methods within computer graphics have made progress on related problems under the framing of inverse procedural modeling. These methods aim to infer structured symbolic objects that explain visual inputs, e.g. procedural models that produce visual outputs when executed. A typical framing these approaches take is to induce a grammar with a bottom-up procedure, e.g. Bayesian merging (Hwang et al., 2011). These techniques have demonstrated success across many visual domains, including plants (Stava et al., 2010; Guo et al., 2020) and buildings (Nishida et al., 2016; Martinovic & Van Gool, 2013; Nishida et al., 2018; Demir et al., 2016), and some can even induce more general probabilistic programs (Ritchie et al., 2018). Unfortunately, these methods lack the generality we desire: they are not able to induce grammars outside of their specialized domains and often require structured input data. In contrast, our approach uses learned networks that guide a top-down inference procedure.
Relatedly, another line of work has explored how to infer procedural models that explain a single visual input; we refer to this task as visual program induction (VPI) (Ritchie et al., 2023). Though some VPI methods are non-learning based, relying on search and heuristics (Du et al., 2018; Xu et al., 2021), recent methods have investigated learning-based solutions for this task. While many of these VPI approaches are specialized for domains of interest such as 3D CAD modeling (Xu et al., 2022; Tian et al., 2019), others have proposed learning techniques that work over multiple domains (Jones et al., 2022; Ellis et al., 2019; Ganin et al., 2018; Hewitt et al., ; Ellis et al., 2021). Similar to these approaches, we aim to learn networks that can infer programmatic objects across domains, but instead of inferring a single program that explains one input, we aim to infer a Template Program that captures a group of visual inputs.
@@ 4.1 Visual Domains
Omniglot Characters  Lake et al. (2015) introduced the Omniglot dataset which contains handwritten characters from 50 languages. These characters are split between a background set (964 characters) and a generalization set (659 characters), where each concept comes with 20 examples. We use the background characters for training and validation sets and test on the generalization characters. Our DSL for drawing characters produces strokes by moving a virtual pen. The pen moves at an angle, for varying distances, optionally bowing inwards or outwards. It can be lifted up or put down and has the option to back-track to previous positions. As we are more interested in modeling stroke structure than physical handwriting dynamics, we adopt a simplified ink model compared with previous work: any pixel the pen passes through is filled completely.
@@
Method FDnormal-$\Downarrow$ Confnormal-$\Uparrow$ MMDnormal-$\Downarrow$ Covnormal-$\Uparrow$ mIoUnormal-$\Uparrow$ FDnormal-$\Downarrow$ Confnormal-$\Uparrow$ MMDnormal-$\Downarrow$ Covnormal-$\Uparrow$ mIoUnormal-$\Uparrow$ FDnormal-$\Downarrow$ MMDnormal-$\Downarrow$ Covnormal-$\Uparrow$ mIoUnormal-$\Uparrow$
@@ 4.3 Concept Few-shot generation
We present quantitative few-shot generation results in Table 1 (details in Appendix F). For each domain and test-set concept, we provide every method with a group of 5 visual inputs and ask it to synthesize 5 generations. Comparing these generations to a reference set that contains held-out examples from the same concept, we compute the following metrics using the latent space of a domain-specific auto-encoder: Frechet Distance (FD), Minimum Matching Distance (MMD), and Coverage (Cov). For the Omniglot and 2D layout domains, we also report class confidence (Conf), the average predicted probability of each generation being a member of the target class under a classifier trained on all domain concepts.
@@ 4.4 Concept Co-segmentation
We compare how our method does on co-segmentation tasks across domains. Our main comparison is against BAE-NET (Chen et al., 2019), which is designed specifically for this task. For Omniglot, BPL and GNS can also perform this task by parsing visual inputs to ordered strokes. We report results of our experiments in Table 1. We evaluate performance with a mean intersection over union metric (mIoU) that measures how closely the output segmentation predictions match the ground-truth labelings. Despite the fact that our method never trains on human stroke data, we achieve a better mIoU on this co-segmentation task compared with GNS, and nearly match the metric value achieved by BPL. Though our output co-segmentations are less structured compared with the ordered stroke parses BPL and GNS can produce, we are encouraged by our method’s performance in this task. For our 3D shapes domain, as BAE-NET was originally designed to operate over voxels, our comparisons against it use a variant of our method that also takes in voxel inputs. We visualize an example co-segmentation of each method in Fig. 4. Across domains and input modalities, we find that we outperform BAE-NET for this task.
@@ 4.5 Discussion
Though we mainly evaluate our method on few-shot generation and co-segmentation, these are not the only concept-related tasks our framework can support. For Omniglot, we explore how our approach can be used for unconstrained concept generation. In fact, this is a task we naturally solve as part of our fine-tuning procedure: the wake-sleep component of each training loop uses an unconditional generative model to sample Template Programs that represent new concepts. We visualize some of these generations in Fig. 5.
Table 2: We run a perceptual study to evaluate few-shot generation, finding our method is greatly preferred over task-specific alternatives and slightly preferred over domain-specific alternatives.
@@ 5 Conclusion
There are a number of directions we would like to investigate in future work. So far, we have only considered visual input groups of size 5, but this constraint can be relaxed by changing the conditioning information presented to the TemplateNet. More generally, one could train this network over input groups of varying sizes (i.e through random masking of visual encodings), to achieve the most inference time flexibility. Further, while we allow our Template Programs to capture parametric relations, the kinds of relations we have so far investigated are fairly simple: static variable assignment and variable re-use. For continuous variables especially, it would be interesting to consider learning more complex relations. These could be declarative (e.g. closed-formula equations that operate over other parameters) or could even describe distributions (e.g. by parameterizing a Gaussian). Our Template Programs framework offers a promising step forward towards general concept learning, and looking forward, we are excited to see how it can be adapted to an even broader array of tasks and domains.
Broader Impact
@@ Appendix A Appendix Overview
We overview the contents of our appendices and supplemental material. In Section B we provide additional experimental results. In Section C we provide more information concerning our various visual domains. In Section D we provide details of our learned models. In Section E we provide details on how we design our training procedure. In Section F we provide further details of our experimental design. In Section G we describe implementation considerations of each alternative we compare our system with. Finally, in Section H we provide pseudo-code for select parts of our method.
In our accompanying supplemental material we additionally provide (i) qualitative few-shot generation results, (ii) qualitative co-segmentation results, (iii) all images used in our perceptual study, (iv) further information on our procedurally generated target dataset for the layout domain (including visualizations of every concept and details on our train-val-test split), and (v) visualizations of synthetic training data from each domain.
@@ B.2 Method Ablation Study
We run an ablation study to validate different design decisions of our method. We compare our described system against the following variants. Ours - rel is a variant of our method where we remove parameter relationships from Template Programs. As by default we only support parameter relationships for argument types that take on discrete values (i.e. categorical variables) we also investigate a variant of our system that adds parameter relations (static assignment and reuse) for float-typed arguments: Ours + float rels. We also compare against a version of our method where we remove HOLE tokens, so that instantiations from Template Programs always use the same function call sequence: Ours - HOLE. Next we compare against a variant where we remove the Structural Expansion step, so the ParamNet must produce a program from the Template Program directly. As it doesn’t see the $SE$ intermediary result, it must fill in HOLE tokens while figuring out how to predict parameter values. We call this variant Ours - $SE$. Finally, we compare against a variant of our base method without any finetuning, where networks only get to train on synthetic data: Ours - finetune
@@ Appendix C Domain Details
For the 3D shape domain, we additionally provide details on how we produce our target dataset. We provide additional details concerning our procedurally generated layout domain in the supplemental material. While we have previously explained how we divide concepts between training and test sets for each of our domains, we have not yet mentioned how we divide training examples into a validation set. We find that a simple approach of taking a subset of training concepts with fixed exemplars as a ‘validation’ set works well in practice. This validation set controls different early stopping components of our finetuning procedure, but otherwise these concepts are not given special treatment (i.e. they are not removed from the finetuning training set).

### v2

@@ Abstract
neurosymbolic methods, visual programs, concept learning, program induction, inverse graphics, procedural modelings
@@ 1 Introduction
We experimentally validate that our method is capable of inferring Template Programs across multiple visual domains: 2D layouts, Omniglot characters, and 3D shapes. We demonstrate that Template Programs natively support a number of downstream applications, including few-shot generation and co-segmentation. We are unaware of any other method that is able to perform these tasks in a domain-general fashion, so we compare against either task-specific or domain-specific alternatives. With respect to task-specific approaches, we find that our neurosymbolic method achieves superior performance. For the one domain, Omniglot (Lake et al., 2015), where task-general methods have been proposed, we compare our domain-general method against domain-specific approaches and find that we are able to achieve competitive performance. We release code for our experiments at: https://github.com/rkjones4/TemplatePrograms
@@ 2 Related Work
Concept Learning Many works have studied concept learning and related tasks (Lake et al., 2019b). A number of methods take a task-specific focus, such as training neural networks to perform few-shot classification (Vinyals et al., 2016; Snell et al., 2017). Other approaches have focused instead on generation, under different conditioning frameworks (Rezende et al., 2016; Giannone et al., 2022). While these methods perform well in the areas they specialize for, none of them achieve the task-flexibility we desire.
@@
Inverse Procedural Modeling While not often presented as concept learners, methods within computer graphics have made progress on related problems under the framing of inverse procedural modeling. These methods aim to infer structured symbolic objects that explain visual inputs, e.g. procedural models that produce visual outputs when executed. A typical framing these approaches take is to induce a grammar with a bottom-up procedure, e.g. Bayesian merging (Hwang et al., 2011). These techniques have demonstrated success across many visual domains, including plants (Stava et al., 2010; Guo et al., 2020) and buildings (Nishida et al., 2016; Martinovic & Van Gool, 2013; Nishida et al., 2018; Demir et al., 2016), and some can even induce more general probabilistic programs (Ritchie et al., 2018). Unfortunately, these methods lack the generality we desire: they are not able to induce grammars outside of their specialized domains and often require structured input data. In contrast, our approach uses learned networks that guide a top-down inference procedure.
Relatedly, another line of work has explored how to infer procedural models that explain a single visual input; we refer to this task as visual program induction (VPI) (Ritchie et al., 2023). Though some VPI methods are non-learning based, relying on search and heuristics (Du et al., 2018; Xu et al., 2021), recent methods have investigated learning-based solutions for this task. While many of these VPI approaches are specialized for domains of interest such as 3D CAD modeling (Xu et al., 2022; Tian et al., 2019), others have proposed learning techniques that work over multiple domains (Jones et al., 2022; Ellis et al., 2019; Ganin et al., 2018; Hewitt et al., ; Ganeshan et al., 2023). Similar to these approaches, we aim to learn networks that can infer programmatic objects across domains, but instead of inferring a single program that explains one input, we aim to infer a Template Program that captures a group of visual inputs.
A few works have considered how partial program constructs can aid in program synthesis. Sketch uses partial programs to constrain a SAT-based inductive synthesis procedure (Solar-Lezama, 2008). While this framing is compelling for discrete domains, adapting this technique for more complex visual domains often requires first solving a primitive decomposition problem (Ellis et al., 2018). DreamCoder is a system that takes as input a corpus of tasks and a base DSL, and from this uses an iterative bootstrapping procedure to find a library of abstraction functions (Ellis et al., 2021). While these discovered abstractions may contain holes, they cannot be used to perform concept-related tasks. In contrast to both of these works, which use partial program intermediates to find a deterministic program that matches a single input, our method aims to find a partial programs that explains a collection of visual inputs.
@@ 4.1 Visual Domains
Omniglot Characters  Lake et al. (2015) introduced the Omniglot dataset which contains handwritten characters from 50 languages. These characters are split between a background set (964 characters) and a generalization set (659 characters), where each concept comes with 20 examples. We use the background characters for training and validation, and test on the generalization characters. Our DSL for drawing characters produces strokes by moving a virtual pen. The pen moves at an angle, for varying distances, optionally bowing inwards or outwards. It can be lifted up or put down and has the option to back-track to previous positions. As we are more interested in modeling stroke structure than physical handwriting dynamics, we adopt a simplified ink model compared with previous work: any pixel the pen passes through is filled completely.
@@
[…]
@@ 4.3 Concept Few-shot generation
We present quantitative few-shot generation results in Table 1 (details in App. F). For each domain and test-set concept, we provide every method with a group of 5 visual inputs and ask it to synthesize 5 generations. Comparing these generations to a reference set of held-out examples from the same concept, we compute the following metrics using the latent space of a domain-specific auto-encoder: Frechet Distance (FD), Minimum Matching Distance (MMD), and Coverage (Cov). For the Omniglot and 2D layout domains, we also report class confidence (Conf), the average predicted probability of each generation being a member of the target class under a classifier trained on all domain concepts.
@@ 4.4 Concept Co-segmentation
We compare how our method does on co-segmentation tasks across domains. Our main comparison is against BAE-NET (Chen et al., 2019), which is designed specifically for this task. For Omniglot, BPL and GNS can also perform this task by parsing visual inputs to ordered strokes. We report results of our experiments in Table 1. We evaluate performance with a mean intersection over union metric (mIoU) that measures how closely the output segmentation predictions match the ground-truth labelings. Despite the fact that our method never trains on human stroke data, we achieve a better mIoU on this co-segmentation task compared with GNS, and nearly match the metric value achieved by BPL. Though our output co-segmentations are less structured compared with the ordered stroke parses BPL and GNS can produce, we are encouraged by our method’s performance in this task. For our 3D shapes domain, as BAE-NET was originally designed to operate over voxels, our comparisons against it use a variant of our method that also takes in voxel inputs. We visualize an example co-segmentation of each method in Figure 4. Across domains and input modalities, we find that we outperform BAE-NET for this task.
@@ 4.5 Discussion
Though we mainly evaluate our method on few-shot generation and co-segmentation, these are not the only concept-related tasks our framework can support. For Omniglot, we explore how our approach can be used for unconditional concept generation. In fact, this is a task we naturally solve as part of our fine-tuning procedure: the wake-sleep component of each training loop uses an unconditional generative model to sample Template Programs that represent new concepts. We visualize some of these generations in Figure 5.
Table 2: Perceptual study results evaluating few-shot generation performance. Our method is greatly preferred over task-specific alternatives and slightly preferred over domain-specific alternatives.
@@ 5 Conclusion
There are a number of directions we would like to investigate in future work. So far, we have only considered visual input groups of size 5, but this constraint can be relaxed by changing the conditioning information presented to the TemplateNet. More generally, one could train this network over input groups of varying sizes (i.e through random masking of visual encodings), to achieve the most inference time flexibility. Further, while we allow our Template Programs to capture parametric relations, the kinds of relations we have so far investigated are fairly simple: static variable assignment and variable re-use. For continuous variables especially, it would be interesting to consider learning more complex relations. These could be declarative (e.g. closed-formula equations that operate over other parameters) or could even describe distributions (e.g. by parameterizing a Gaussian). Our Template Programs framework offers a promising step forward towards general concept learning, and looking ahead, we are excited to see how it can be adapted to an even broader array of tasks and domains.
Acknowledgments
We would like to thank the anonymous reviewers for their helpful suggestions and all of our perceptual study particpants for their time. Renderings of 3D shapes were produced using the Blender Cycles renderer. This work was funded in parts by NSF award #1941808 and a Brown University Presidential Fellowship. Daniel Ritchie is an advisor to Geopipe and owns equity in the company. Geopipe is a start-up that is developing 3D technology to build immersive virtual copies of the real world with applications in various fields, including games and architecture. Part of this work was done while R. Kenny Jones was an intern at Adobe Research.
Impact Statement
@@ Appendix A Appendix Overview
We overview the contents of our appendices. In Section B we provide additional experimental results. In Section C we provide more information concerning our various visual domains. In Section D we provide details of our learned models. In Section E we provide details on how we design our training procedure. In Section F we provide further details of our experimental design. Finally, in Section G we describe implementation considerations of each alternative we compare our system with.
@@ B.2 Method Ablation Study
We run an ablation study to validate different design decisions of our method. We compare our described system against the following variants. Ours - rel is a variant of our method where we remove parameter relationships from Template Programs. As by default we only support parameter relationships for argument types that take on discrete values (i.e. categorical variables) we also investigate a variant of our system that adds parameter relations (static assignment and reuse) for float-typed arguments: Ours + float rels. We also compare against a version of our method where we remove HOLE tokens, so that instantiations from Template Programs always use the same function call sequence: Ours - HOLE. Here, we task our network to specify a single program structure that is applicable across the group without using HOLE tokens, and it is still responsible for declaring parameter relationships. As there are no HOLE tokens, the ExpansionNet will not be used, but the ParamNet will still be used to figure out how the instantiations of the Template Program should be parameterized. Next we compare against a variant where we remove the Structural Expansion step, so the ParamNet must produce a program from the Template Program directly. As it doesn’t see the $SE$ intermediary result, it must fill in HOLE tokens while figuring out how to predict parameter values. We call this variant Ours - $SE$. Finally, we compare against a variant of our base method without any finetuning, where networks only get to train on synthetic data: Ours - finetune
@@ B.3 Unconditional Concept Generation
B.4 Visual Concept Groupings
Typically, past concept learning approaches have assumed access to a dataset that is structured according to visual concepts. For instance, systems like VHE or FSDM require the ability to sample groups of input from the same visual concept during training. This is the same amount of dataset structure that our method requires: during fine-tuning we randomly sample “tasks” according to these visual concept groupings. Note that this requirement is less stringent than many inverse procedural modeling systems, and the BPL and GNS systems, that additionally require per-object structural annotations.
The Omniglot dataset was designed with this kind of visual concept decomposition in mind: each example data-point corresponds with exactly one character type. We design our layout domain in the spirit of Omniglot: each image in the layout domain is associated with a single concept. Following past work, on these domains we always assume “valid” input groups, such that each member is from the same visual concept.
However, this type of clean partition is not as easy to find for 3D shape structures. As there are no known datasets that group shape structures into visual concepts, we propose a heuristic method for forming approximate visual concepts out of shape structures (Appendix  C.3). The concept groups we find under this formulation have different levels of consistency among their members (where we say a less consistent group forms a “harder” input problem).
For instance consider the examples shown in Figure 2. A chair with a regular base and vertical slats (row 7, col 6) could be in one group with only chairs that also have regular bases and vertical slats (row 7, col 7) or (like in the example we show) could also be grouped with chairs that have backs with horizontal slats (row 7, col 8). In our paradigm, the group of visual inputs (along with our objective function) implicitly defines the granularity of the target visual concept. In this case, the latter grouping is considerably harder to handle for concept learning tasks, as it requires a method that is able to reason over input groups that partially mismatch on structures.
Our Template Programs framework is capable of handling even difficult input groups; our partial program formulation allows our system to explicitly maintain the shared structural aspects of the group while leaving HOLE tokens as responsible for representing the aspects of the input group that structurally differ. This design allows us to successfully capture the visual concepts of the two chair groups in Figure 2. The left chair group has filled in chair backs, arm-rests, but alternates between regular and pedestal chair bases. The right chair group has regular leg bases, no arm-rests, but differs between chair backs with horizontal slats or vertical slats. As can be seen in the “gen” row, our system is capable of synthesizing novel shape structures that accord with the structural specifications implied by the input visual groups.
B.5 Reconstruction Performance
Our system learns how to amortize the difficult inverse search problem of finding a Template Program and instantiations that correspond with a group of visual inputs. This search (our inference procedure) is guided by our networks which are trained on a “training corpus” of visual concepts, separate from those we evaluate on.
The “seg” rows in Figure 2 visualize the reconstructions (of the inputs on the top rows) that our method produces. While these reconstructions do not exactly recreate the input, they usually create very good approximations. If reconstruction was our primary goal, it might even be possible to improve the fit through a differentiable execution and refinement procedure.
To explore this phenomenon further, we provide the following reconstruction performance results across our domains in Table 4. We report the reconstruction fit for both the training set and test set visual concepts. To show the benefits of our learning methodology we compare the reconstruction fit from the pretrained version of our networks (that learn only on synthetic data) to the finetuned versions of our networks (that finetune on visual concepts from training set). The metrics we use are (full descriptions in Appendix C):
  • 2D Layout: color-based IoU (higher is better)
  • Omniglot: edge-based chamfer distance (lower is better)
  • 3D Shapes (primitive input): structural corner distance (lower is better)
  • 3D Shapes (voxel input): IoU (higher is better)
Table 4: Comparing reconstruction performance across domains, concept sets, and model versions.
Domain Mode Train Recon Test Recon Test Recon (long)
2D Layout $\uparrow$ Pretrain .822 .808
Finetune .972 .909 .937
Omniglot $\downarrow$ Pretrain .658 .648
Finetune .468 .503 .405
3D Shapes (prim) $\downarrow$ Pretrain .26 .305
Finetune .05 .06 .05
3D Shapes (voxel) $\uparrow$ Pretrain .601 .589
Finetune .865 .83 .851
As demonstrated, our solution is effective at solving this inverse visual program induction problem. For both the training concepts and the held-out test concepts, our finetuning procedure meaningfully improves the reconstruction performance in all cases. For our downstream concept-related tasks we use a more expensive inference procedure (“long” - e.g. increase the beam size, Section 3.2) and this gives even better reconstruction results for test-set concepts (see the numbers in the rightmost column of the table).
While our system offers strong reconstruction performance, it is likely that alternative methods could be used to infer single visual programs that better reconstruct an individual visual input. In contrast, our system learns to solve this visual program induction problem over a group of inputs by going through a shared structural intermediary (a Template Program), which allows us to perform concept-related tasks like few-shot generation and co-segmentation (which prior single instance VPI approaches are not suited for).
B.6 Failure Modes
Recon Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Gen Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: When our method fails to find good reconstructions of an input concept, downstream task performance worsens.
Bad reconstruction
A possible failure mode is that our inference networks can’t find a Template Program whose instantiations well-capture an input visual group with respect to our objective function. In such cases, the few-shot generation and co-segmentation results of our method are typically worse. For instance, consider Figure 7. For two Omniglot examples, in the top row we show the input concept groups, in the middle rows we show the reconstructions from our method, and in the bottom rows we show the few-shot generations from our method. Because the same Template Program is used in both the reconstruction and few-shot generation step, failure in one place often means failure in the other. While from one perspective this is a limitation, a positive view of this phenomena is that our method can provide insight into cases where it is “unsure” about its parse. For instance, it could use the objective function score of its reconstruction as a measure of its confidence on how well it will perform on downstream tasks. Moreover, as we show in Table 4, reconstruction performance can improved by spending more time on inference, which can help to avoid this limitation.
Bad Input Groups
How would our system handle ‘bad’ input groups that contain outliers, or have no commonality among their members? The job of the template network is to consume a group of visual inputs and infer a Template Program that captures the common structure among all members. In such an adversarial setting, it is possible (depending on random sampling) that there are no elements of structure common to all members of the input grouping.
In this case, the “best” result of our system would be to return a “dummy” Template Program that consists of a single HOLE token; this HOLE token would be able to be expanded into any arbitrary z to explain each individual group member.
For typical visual concept groupings, this degenerate solution is discouraged by our objective function, which penalizes description length differences between ”full” programs and their corresponding Template Programs. While finetuning our system with reasonable concept groups we have never observed the system falling-back to this degenerate solution.
Exploring how to extend our framework to handle “noisy” input groupings would be a very interesting direction for future work. This could potentially be approached by (i) extending our objective function to account for outliers (if we want to ignore the distractors) or (ii) adding control flow operators into the DSLs we learn over, which would give the Template Programs an opportunity to account for structural differences without relying solely on HOLE tokens.
@@ Appendix C Domain Details
For the 3D shape domain, we additionally provide details on how we produce our target dataset. While we have previously explained how we divide concepts between training and test sets for each of our domains, we have not yet mentioned how we divide training examples into a validation set. We find that a simple approach of taking a subset of training concepts with fixed exemplars as a ‘validation’ set works well in practice. This validation set controls different early stopping components of our finetuning procedure, but otherwise these concepts are not given special treatment (i.e. they are not removed from the finetuning training set).